OpenStudy (anonymous):

find the general solution of the following differential equation: y dy/dx=sinx

OpenStudy (saifoo.khan):

-cos x

OpenStudy (anonymous):

what are the steps i dont just need the answer...

OpenStudy (anonymous):

please...

OpenStudy (saifoo.khan):

I just learned like that. im not sure if there's any working for this. sorry.

OpenStudy (anonymous):

Multiply both sides by dx to get: \[ ydy = sinx dx \]Integrate both sides: \[ \frac{y^2}{2} = -cos(x) + C_1 \implies y = \pm \sqrt{C - 2cosx}\]

OpenStudy (anonymous):

thxx

OpenStudy (anonymous):

You are welcome :-)